OpenStudy (anonymous):

the radius of a circle is decreasing at a constant rate of .1 centimeter per second. in terms of circumference C, what is the rate of change of the area of the circle, in square centimeters per second?

OpenStudy (anonymous):

yup. then mukltiply by whatever coefficents there are for your 4 numbers

OpenStudy (anonymous):

A = pi*r^2; dA/dt = 2*pi*r*dr/dt. We know dr/dt = 0.1cm/sec. We need to know the 'r' at which we want to determine the change in area (dA/dt) and stick that into the equation: dA/dt = 2*pi*r*dr/dt

OpenStudy (anonymous):

wait, why do you multiply dr/dt

OpenStudy (anonymous):

seocnd statement after Area formula

OpenStudy (anonymous):

This is a related rate problem. The dA/dt is related to the dr/dt, which is a given.

OpenStudy (anonymous):

Yes, you want dA/dt, which equals 2*pi*r*dr/dt

OpenStudy (anonymous):

so.. the anser is -.1(2pirR)

OpenStudy (anonymous):

WAIT NO